OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 5 2x - 4y = 10 Infinitely many solutions No solutions (0, 0) (5, 10)

OpenStudy (igreen):

x - 2y = 5 Add 2y to both sides: x = 2y + 5 Substitute '2y + 5' for 'x' in the 2nd equation: 2x - 4y = 10 2(2y + 5) - 4y = 10 Distribute 2 into the parenthesis: 4y + 10 - 4y = 10 Combine like terms, what's 4y - 4y?

OpenStudy (anonymous):